4 over the square root of 2x
Answer by ilana(307)

A fraction with a radical should not have a radical in the denominator- this is not considered simplified. But when you square a square root, you get an integer. You can always multiply something by 1. (square root of 2x)/(square root of 2x)=1. So multiply (4/sqrt2x)*(sqrt2x/sqrt2x)=(4*sqrt2x)/(sqrt2x*sqrt2x)= 4sqrt2x/(sqrt2x)^2= 4sqrt2x/2x= (2sqrt2x)/x.